OpenStudy (anonymous):

A ramp 21 feet in length rises to a loading platform that is 17 feet off the ground. What is the angle of elevation of the ramp? Answer in units of◦

OpenStudy (anonymous):



OpenStudy (anonymous):

sin(theta)=17/21

OpenStudy (anonymous):

SOH = sin (opposite over hypotenuse so when dealing with trig function in order to get that cos on the other side multiply by inverse so your answer will be \[\theta=\arcsin(\frac{17}{21})\]

OpenStudy (anonymous):

sin on the other side* so your answer will be approximately 54 degrres